[Brahms-l] Shift report 20050320 16:00-24:00

From: <apache@brahms-web.brahms.bnl.gov>
Date: Sun Mar 20 2005 - 00:04:19 EST
Shifter leader was sjs
Shifters were 


Good beam throughout shift. Lost 1/2 hour for new store, but otherwise continuous running. Finished 1/2A;6B field settings and have changed to 1/2B; 6A setting.
For 1/2A;6B:
           trig2    trig3
run 13870   4.8K    148.6K
    13871   4.5K    133.4K
    13872   3.3K     96.5K
For 1/2B;6A:
           trig2    trig3
run 13873   0.9K     16.3K
    13876   7.5K    164.1K
    13877   6.7K    138.9K
    13878   5.6K    112.1K

started run 13879 at end of shift...next beam dump scheduled for 1:35 am.

No new detector problems noticed, although gain of C4 ADC9&10 seems too low.

_______________________________________________
Brahms-l mailing list
Brahms-l@lists.bnl.gov
http://lists.bnl.gov/mailman/listinfo/brahms-l
Received on Sun Mar 20 00:04:36 2005

This archive was generated by hypermail 2.1.8 : Sun Mar 20 2005 - 00:04:49 EST